OpenStudy (anonymous):

Solve the system. y = -1/3 x + 2 and x + 3y = 3 (0, 1) (1, 0) (3, 1/3) (3/2, 0) no solution

OpenStudy (goformit100):

no solution Simple Its because solve the equations or Put the Values in it ...

OpenStudy (anonymous):

Making x=-3y+3 and plugging the first equation of the system. y=-1/3(3-3y)+2 y=-1+y+2; y-y=-1+2; 0=1. This is false.